In the arrangement shown , assume negligible friction b/w blocks and table. Show that the tension in the string is 2 m1F / (4m1+m2)

4403-2318_3899_laws2.bmp


In the arrangement shown , assume negligible friction b/w blocks and table. Show that the tension in the string is 2 m1F / (4m1+m2)

Grade:12

1 Answers

Pratham Ashish
14 years ago

hi,

 see if the block m2 moves   x distance due to force F , then by the constranit of constant length of string  m1 has to move 2x distance

 from here we can conclude that ,

  accl. of m1 = 2 accl. of m2

 a1 = 2 a2  ....................(1)

for block 1 ,

T = m1 a1.........

or, T = m1 *2 * a2........(2)

 for block 2,

 F - 2T = m2 a2 ..................(3)

from Eq. 2 &3 ,elliminating T, we get

F = ( m2 + 4 m1 ) a2   , putting in (2)

  T = 2m1 F / (4m1 + m2 )